• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name C. C. Chang

    Log entry time 16:03:50 on November12,2002

    Entry number 89971

    keyword=Shift Summary

    Run# Start Stop Event# Comments
    3115 08:04 09:15 1M
    3116 09:16 09:54 1M
    3117 09:57 10:29 1M
    3118 10:49 11:24 1M
    3119 11:25 12:05 1M

    08:55 MCC called to inform us beam will be off for about 5 minutes.
    09:05 Beam is back.
    09:17 MCC called to tell us that beam will gradually be increased to 100 uA.
    09:25 We have 100 uA of beam; but the beam is not stable.
    10:35 Reboot target IOC.
    12:05 MCC called to let us know that they are taking beam away for
    about two hours for tuning.
    14:00 MCC called to let us know that beam will be resumed.
    14:45 MCC called to try to send CW beam to the HALL. They are still tuning
    the beam, so it is not ready for us to resume data taking as yet.
    16:00 Still no stable as of now...